OpenStudy (anonymous):

What is the solution to the following system of equations? y = -x2 - 5x - 4 y = -x2 + 9x - 18

OpenStudy (anonymous):

is it -2x or -x^2

OpenStudy (anonymous):

-x^2

OpenStudy (anonymous):

-x^2-5x-4=-x^2+9x-18 Can you solve that?

OpenStudy (anonymous):

im not sure haha could you help me

OpenStudy (anonymous):

\[-x^2 -5x-4=-x^2 +9x-18\] first add x^2 to both sides (-5x-4=9x-18) Then add 5x to both sides (-4=14x-18) Then add 18 to both sides (14=14x) Then divide 14x from both sides (x=1) Now subsitute x in the equation and see what y equals